OpenStudy (anonymous):

Find the period of y = -5cosx

OpenStudy (campbell_st):

the general form is y = acos(bx) a is the amplitude and the period is \[period = \frac{2\pi}{b}\] so what is the coefficient of x, when you find it... that will be b hope it helps

OpenStudy (anonymous):

Perfect, thanks.